OpenStudy (sammietaygreen):

if (fx) = x^2 + 1 and g(x) = 2x, find [f*g](x) how do I solve this??

OpenStudy (jdoe0001):

\(\bf f(x)= x^2 + 1 \qquad g(x) = 2x \\ \quad \\\ [f\cdot g](x)\implies f(x)\cdot g(x)\implies (x^2+1)\cdot (2x)\implies ?\)

OpenStudy (sammietaygreen):

Would I just factor it??

OpenStudy (jdoe0001):

well, is just a multiplication

OpenStudy (sammietaygreen):

2x(x^2+1)?

OpenStudy (jdoe0001):

yes

OpenStudy (sammietaygreen):

not one of the answer choices :// 2x^2+1 is one of them